OpenStudy (anonymous):

a 4 hr river cruise goes 20 km upstream and then back again. the river has a current of 3 km an hr going downstream. what is the boat's speed in still water and how long was the upstream journey in minutes? round to the nearest tenth for each

OpenStudy (michele_laino):

I think that, during the upstream journey, the relative speed boat with respect to the riverbank is v+3, where v is the speed of the boat with respect to the water. so time employed by the boat to make the upstream journey, is: \[\frac{ d }{ v+3 }\] where d= 20Km

OpenStudy (michele_laino):

during the downstream journey, our boat has relative speed v-3, with respect to the riverbank, so the time employed by our boat to make the downstream journey, is: \[\frac{ d }{ v-3 }\]

OpenStudy (michele_laino):

now, if we add both times each other, we can get this: \[\frac{ d }{ v+3 }+\frac{ d }{ v-3 }=4\] substituting for d, we have: \[\frac{ 20 }{ v+3 }+\frac{ 20 }{ v-3 }=4\]

OpenStudy (michele_laino):

please note that the above is a quadratic equation. Solving for v, we get two solutions, even if only one of them is acceptable. Our acceptable solution is: \[v=5+\sqrt{34}=10.8 Km/hour\]

OpenStudy (michele_laino):

so time employed by the boat for upstream journey, is: \[\frac{ 20 }{ 10.8+3 }=...\] please, complete the computation

OpenStudy (michele_laino):

if something is not clear, please tell me!

OpenStudy (anonymous):

1.5 hrs?

OpenStudy (michele_laino):

better is 1.4 hrs